# Injecting Syntactic Sub-Networks in Transformer Language Models with Tree Regularization [Uncaptioned image]

### v1

@@ Abstract
While compositional accounts of human language understanding are based on a hierarchical tree-like process, neural models like transformers lack a direct inductive bias for such tree structures. Introducing syntactic inductive biases could unlock more robust and data-efficient learning in transformer language models (LMs), but existing methods for incorporating such structure greatly restrict models, either limiting their expressivity or increasing inference complexity. This work instead aims to softly inject syntactic inductive biases into given transformer circuits, through a structured regularizer. We introduce TreeReg, an auxiliary loss function that converts bracketing decisions from silver parses into a set of differentiable orthogonality constraints on vector hidden states. TreeReg integrates seamlessly with the standard LM objective, requiring no architectural changes. LMs pre-trained with TreeReg on natural language corpora such as WikiText-103 achieve up to 10% lower perplexities on out-of-distribution data and up to 9.5 point improvements in syntactic generalization, requiring less than half the training data to outperform standard LMs. TreeReg still provides gains for pre-trained LLMs: Continued pre-training of Sheared Llama with TreeReg results in improved syntactic generalization, and fine-tuning on MultiNLI with TreeReg mitigates degradation of performance on adversarial NLI benchmarks by 41.2 points.
Sneaking Syntax into Transformer Language Models with Tree Regularization [Uncaptioned image]
@@ 1 Introduction
A substantial body of research (Crain and Nakayama, 1987; Pallier et al., 2011; van Schijndel et al., 2013; Hale et al., 2018) suggests that human language processing is inherently hierarchical: syntactic rules combine word-level meanings to give rise to semantics for larger constituents, which are then combined into sentences. In contrast, transformers (Vaswani, 2017), the architecture behind large language models (LLMs), allow for arbitrary, non-hierarchical routing of information while processing an input sentence.
LLMs are pre-trained on massive amounts of data to achieve reasonable generalization Achiam et al. (2023); Dubey et al. (2024), but recent studies find that state-of-the-art LLMs still struggle with compositional generalization Guo et al. (2020); Berglund et al. (2023), the ability to understand familiar words in novel contexts. Prior work Socher et al. (2013); Eriguchi et al. (2016) proposed explicitly tree-structured models capable of data-efficient compositional generalization, but they are too constrained and not amenable to pre-training at scale.
@@
The use of TreeReg improves syntactic generalization across model and data scales. On diagnostic sentence transformation tasks such as tense inflection and question formation (McCoy et al., 2020), it significantly enhances grokking behavior (§ 4). It also improves syntactic generalization on the BLiMP (Warstadt et al., 2020) and SyntaxGym Gauthier et al. (2020) test suites by up to 3.2 and 9.5 points, both when used during pre-training from scratch (§ 5), and during continued pre-training (§ 6.1). LMs with TreeReg surpass the syntactic generalization of baseline LMs with less than half the training data (§ 5.2). Most importantly, TreeReg improves out-of-distribution language understanding. When used during pre-training of LMs on fully parsed (§ 5.1) as well as a mixture of parsed and unparsed (§ 5.3) data, TreeReg reduces out-of-distribution perplexities by up to 9.2%. When TreeReg is used for fine-tuning an LLM on MultiNLI (Williams et al., 2018), it mitigates performance drops by up to 41.2 points on adversarial NLI benchmarks (§ 6.2).
@@ 2 Background and Notation
Consider a sentence $S=\{x_{1},x_{2},\ldots,x_{n}\}$ where $x_{j}$ is the $j$${}^{\hbox{th}}$ token in the sentence. Let tree $T(S)$ over sentence $S$ be a set of spans $S_{a;b}=\{x_{a},\ldots,x_{b}\}$ such that if $S_{a;b}$ and $S_{b+1;c}\in S$, then $S_{a;c}\in S$. We say $S_{a;c}$ is split at index $b$ to get spans $S_{a;b}$ and $S_{b+1;c}$ of $T(S)$. Prior work (Pallier et al., 2011; van Schijndel et al., 2013; Hale et al., 2018) suggests that human language processing is hierarchical—sentences contain “constituents” or spans of words that act as semantic units, with smaller constituents (“is”, “very happy”) recursively combining to form larger ones (“is very happy”), building up meaning incrementally. The constituency parse $\hat{T}(S)$ is the tree form by all constituents of $S$. Tree-structured models (Socher et al., 2013; Tai et al., 2015) build sentence representations through recursively bottom-up computation over constituency parses, whereas for transformers, any hierarchically structured computation is at most implicitly learned in the attention layers.
@@ 3 Our Approach
TreeReg is a differentiable loss function that injects the hierarchical structure of input constituency parses into transformer circuits. Inspired by Murty et al. (2023c), TreeReg specifies a set of soft constraints to ensure that constituent representations are maximally independent of surrounding context. For any span, we quantify this independence with the Span Contextual Independence Score ($\mathrm{SCIN}{}$, § 3.1). TreeReg maximizes $\mathrm{SCIN}{}$ for spans corresponding to input constituents while simultaneously minimizing it for other spans (§ 3.2). In an optimally trained model, the tree with the highest cumulative $\mathrm{SCIN}{}$ thus recovers the constituency parse. We then provide a greedy algorithm to extract the parse tree induced by TreeReg from a given model for any input sentence.
@@ 3.3 Implementation Details
$\mathrm{SCIN}{}$ computation does not require additional model calls and can be performed during the forward pass that computes $\mathcal{L}_{\text{LM}}{}$. $\mathrm{SCIN}{}$ scores for all spans in a sentence can be calculated simultaneously and efficiently using vectorization, as detailed in Appendix B. Additionally, updates from $\mathcal{L}_{\text{LM}}{}$ and $\mathcal{L}_{\text{TR}}{}$ can occur at different frequencies during training, for example, $\mathcal{L}_{\text{TR}}{}$ may be applied once every $k$ steps of $\mathcal{L}_{\text{LM}}{}$. These factors combine to make TreeReg efficient, leading to an increase of around $25\%$ in training time if $\mathcal{L}_{\text{TR}}{}$ is applied once every 10 $\mathcal{L}_{\text{LM}}{}$ steps, on $25\%$ of attention heads, as per our implementation. This excludes the time required for parsing the data used in TreeReg. The parses are either included with the dataset (BLLIP-LG, MultiNLI) or can be computed once and reused for every training run.
@@ 4 Warm-up: Improving Grokking on Sentence Transformation Tasks
We start by training transformer LMs on two diagnostic tasks derived from PCFGs (See Appendix C for examples and data statistics). In Tense Inflection (TI), the model is provided with an input in the past tense, and required to generate the same input in the present tense. In Question Formation (QF), the model is required to transform a declarative sentence into a question. For QF, we report first word accuracy of the decoded question, and for TI, we report the fraction of test inputs for which the target verb is correctly inflected.
@@ 5.1 Language Modeling on BLLIP-LG
We train 16-layer LMs on the BLLIP-LG dataset (Hu et al., 2020) (Base LM, detailed hyperparameters in Appendix E). For TreeReg LM, we use $\mathcal{L}_{\text{TR}}{}$ at layer 12 of the LM, on 2 out of 8 attention heads, once every 10 steps of $\mathcal{L}_{\text{LM}}{}$, on the already-parsed BLLIP-LG dataset. An exploration of the layer and attention heads used for $\mathcal{L}_{\text{TR}}{}$ can be found in Appendix F. We benchmark syntactic generalization using the BLiMP and SG test suites (See § 2). To evaluate out-of-distribution generalization, we report perplexity on the Penn TreeBank (PTB; Marcus et al., 1993) test set.
Model Syntactic Generalization ($\uparrow$) PPL ($\downarrow$)
BLiMP SG BLLIP PTB
Base LM 72.2 71.9 21.6 49.1
TreeReg LM 74.8 80.0 22.3 44.6
Table 2: Syntactic Generalization on BLiMP and SG test suites, and perplexities (PPL) on BLLIP-LG test set (BLLIP) and PTB test set (PTB) for LMs trained on BLLIP-LG. TreeReg LMs show better syntactic generalization and better generalization to PTB.
@@ Results.
We report results in Table 2. TreeReg achieves a 2.6 point (pt) gain on BLiMP and 8.1 pt overall gain on SG test suites. Figure 2 shows TreeReg improves over the baseline on 4 out of 6 SG test suites, with a notable 17 pt gain on Licensing. TreeReg LM also generalizes better to PTB with 9.2% lower perplexity, with a marginal 0.7 pt perplexity increase on BLLIP-LG.
@@
While syntactic generalization improves with more training data for both the Base and TreeReg LMs, TreeReg LMs outperform the baseline across all settings. Moreover, the performance gap widens as the amount of training data increases. Notably, TreeReg surpasses the baseline’s maximum syntactic generalization performance on SG by 4.4 points when trained on only 50% of the data, demonstrating substantially more sample-efficient syntactic generalization.
@@ 6.2 Fine-tuning on MultiNLI
We finetune Sheared Llama-1.3B (Base LM) on MultiNLI Williams et al. (2018) (training details and evaluation methodology in Appendix D) to get FT LM. For TreeReg LM, we binarize the parses of the premise and hypothesis sentences provided in the dataset and apply $\mathcal{L}_{\text{TR}}{}$ at layer 16 on 6 out of 24 attention heads, every 10 steps of $\mathcal{L}_{\text{LM}}{}$. We report accuracy on the MultiNLI test set, which includes a matched split from the same sources as the training data, and a mismatched split from different sources. We also present results (Table 5) on two diagnostic datasets: MoNLI Geiger et al. (2020) and MED Yanaka et al. (2019).
@@ 7 Analysis
Although TreeReg only implicitly encodes constituency parses in transformer hidden states, we observe that it demonstrates remarkable parsing capabilities. Parses induced by TreeReg align closely with given parses on all datasets, with F1 scores near or above 90. Parsing performance on PTB and 4000 Questions is slightly better with TreeReg CPT LM, likely due to improved representations from Sheared Llama-1.3B (which benefits from large-scale pre-training) on these datasets, that are relatively out-of-distribution from the BLLIP-LG training data. Notably, induced parses from TreeReg align better with groundtruth constituency parses than Benepar (Kitaev et al., 2022) on the 4000 Questions datasets, by 6.8 F1 pts. We provide some example parses on each dataset in Appendix H.
@@
The F1 scores reveal that at layer 1, the parses are nearly random. As we progress to layer 12, they increasingly align with given parses. Beyond layer 12, there is a sharp drop in F1 scores back to near-random by layer 16. This suggests that the circuit involved in $\mathcal{L}_{\text{TR}}{}$ becomes increasingly aligned with given parses up to layer 12, and then rapidly loses this structure. We conclude that to perform next-word prediction, TreeReg LMs implement a non-syntactic function over the syntactic representations built at layer 12. These syntactic representations are, in turn, built gradually over the preceding layers of the transformer, even in the absence of explicit supervision from TreeReg at these layers. In Appendix G, we also show that improvements from TreeReg come specifically from hierarchical computation that align closely with given parses.
@@ 8 Other Related Work
A substantial body of research has pointed out the limitations of transformer-based LMs in achieving robust compositional generalization. They have been shown to struggle with simple tasks, such as determining the parity of binary strings (Bhattamishra et al., 2020; Chiang and Cholak, 2022), balancing bracket sequences (Hahn, 2020), as well as more complex diagnostic natural language tasks (McCoy et al., 2020; Geiger et al., 2020; Dziri et al., 2023). Large-scale pre-training does not fully resolve these issues. For instance, Berglund et al. (2023) shows that GPT-4 (Achiam et al., 2023) fails to generalize on bijective relationships (for example, answering “Who is Tom Cruise’s mother?” correctly but not “Who is Mary Lee South’s son?”).
@@
Building on theories of tree-structured human language processing, prior work explores various tree-structured model architectures (Socher et al., 2013; Tai et al., 2015; Le and Zuidema, 2015; Dyer et al., 2016; Shen et al., 2018; Hu et al., 2021). These models often excel at data-efficient compositional generalization due to their inherently hierarchical computation. However, they have been largely overshadowed by non-hierarchical pre-trained transformer-based LLMs (Brown, 2020; Achiam et al., 2023).

### v2

@@ Abstract
While compositional accounts of human language understanding are based on a hierarchical tree-like process, neural models like transformers lack a direct inductive bias for such tree structures. Introducing syntactic inductive biases could unlock more robust and data-efficient learning in transformer language models (LMs), but existing methods for incorporating such structure greatly restrict models, either limiting their expressivity or increasing inference complexity. This work instead aims to softly inject syntactic inductive biases into given transformer circuits, through a structured regularizer. We introduce TreeReg, an auxiliary loss function that converts bracketing decisions from silver parses into a set of differentiable orthogonality constraints on vector hidden states. TreeReg integrates seamlessly with the standard LM objective, requiring no architectural changes. LMs pre-trained with TreeReg on natural language corpora such as WikiText-103 achieve up to 10% lower perplexities on out-of-distribution data and up to 9.5 point improvements in syntactic generalization, requiring less than half the training data to outperform standard LMs. TreeReg still provides gains for pre-trained LLMs: Continued pre-training of Sheared Llama with TreeReg results in improved syntactic generalization, and fine-tuning on MultiNLI with TreeReg mitigates degradation of performance on adversarial NLI benchmarks by 41.2 points. We release all code to guide future research [^1] https://github.com/ananjan-nandi-9/tree_regularization.
Sneaking Syntax into Transformer Language Models
with Tree Regularization [Uncaptioned image]
@@ 1 Introduction
A substantial body of research (Crain and Nakayama, 1987; Pallier et al., 2011; van Schijndel et al., 2013; Hale et al., 2018) suggests that human language processing is inherently hierarchical: syntactic rules combine word-level meanings to give rise to semantics for larger constituents, which are then combined into sentences. In contrast, transformers (Vaswani et al., 2017), the architecture behind large language models (LLMs), allow for arbitrary, non-hierarchical routing of information while processing an input sentence.
LLMs are pre-trained on massive amounts of data to achieve reasonable generalization Achiam et al. (2023); Dubey et al. (2024), but recent studies find that state-of-the-art LLMs still struggle with compositional generalization Guo et al. (2020); Berglund et al. (2024); Hale and Stanojević (2024), the ability to understand familiar words in novel contexts. Prior work Socher et al. (2013); Eriguchi et al. (2016) proposed explicitly tree-structured models capable of data-efficient compositional generalization, but they are too constrained and not amenable to pre-training at scale.
@@
The use of TreeReg improves syntactic generalization across model and data scales. On diagnostic sentence transformation tasks such as tense inflection and question formation (McCoy et al., 2020), it significantly enhances grokking behavior (§ 4). It also improves syntactic generalization on the BLiMP (Warstadt et al., 2020) and SyntaxGym Gauthier et al. (2020) test suites by up to 3.2 and 9.5 points, both when used during pre-training from scratch (§ 5), and during continued pre-training (§ 6.1). LMs with TreeReg surpass the syntactic generalization of baseline LMs with less than half the training data (§ 5.2). Most importantly, TreeReg improves out-of-distribution language understanding. When used during pre-training of LMs on fully parsed (§ 5.1) as well as a mixture of parsed and unparsed (§ 5.3) data, TreeReg reduces out-of-distribution perplexities by up to 9.2%. When TreeReg is used for fine-tuning an LLM on MultiNLI (Williams et al., 2018), it mitigates performance drops by up to 41.2 points on adversarial NLI benchmarks (§ 6.2). We release all code at https://github.com/ananjan-nandi-9/tree_regularization.
@@ 2 Background and Notation
Consider a sentence $S=\{x_{1},x_{2},\ldots,x_{n}\}$ where $x_{j}$ is the $j$${}^{\hbox{th}}$ token in the sentence. Let tree $T(S)$ over sentence $S$ be a set of spans $S_{a;b}=\{x_{a},\ldots,x_{b}\}$ such that if $S_{a;b}$ and $S_{b+1;c}\in T(S)$, then $S_{a;c}\in T(S)$. We say $S_{a;c}$ is split at index $b$ to get spans $S_{a;b}$ and $S_{b+1;c}$ of $T(S)$. Prior work (Pallier et al., 2011; van Schijndel et al., 2013; Hale et al., 2018) suggests that human language processing is hierarchical—sentences contain “constituents” or spans of words that act as semantic units, with smaller constituents (“is”, “very happy”) recursively combining to form larger ones (“is very happy”), building up meaning incrementally. The constituency parse $\hat{T}(S)$ is the tree form by all constituents of $S$. Tree-structured models (Socher et al., 2013; Tai et al., 2015) build sentence representations through recursively bottom-up computation over constituency parses, whereas for transformers, any hierarchically structured computation is at most implicitly learned in the attention layers.
@@ 3 Our Approach
TreeReg is a differentiable loss function that injects the hierarchical structure of input constituency parses into transformer circuits. Inspired by Murty et al. (2023c), TreeReg specifies a set of soft constraints to ensure that constituent representations are maximally independent of surrounding context. For any span, we quantify this independence with the Span Contextual Independence Score ($\mathrm{SCIN}{}$, § 3.1). TreeReg maximizes $\mathrm{SCIN}{}$ for spans corresponding to input constituents while simultaneously minimizing it for other spans (§ 3.2). In an optimally trained model, the tree with the highest cumulative $\mathrm{SCIN}{}$ thus recovers the constituency parse. We then provide a greedy algorithm to extract the parse tree induced by TreeReg from a given model for any input sentence. A deeper discussion of the design choices underlying our method can be found in Appendix A.
@@ 3.3 Implementation Details
$\mathrm{SCIN}{}$ computation does not require additional model calls and can be performed during the forward pass that computes $\mathcal{L}_{\text{LM}}{}$. $\mathrm{SCIN}{}$ scores for all spans in a sentence can be calculated simultaneously and efficiently using vectorization, as detailed in Appendix C. Additionally, updates from $\mathcal{L}_{\text{LM}}{}$ and $\mathcal{L}_{\text{TR}}{}$ can occur at different frequencies during training, for example, $\mathcal{L}_{\text{TR}}{}$ may be applied once every $k$ steps of $\mathcal{L}_{\text{LM}}{}$. Despite the quadratic time complexity of computing $\mathcal{L}_{\text{TR}}{}$ (which scales with the square of input token count), these factors collectively ensure that TreeReg remains efficient in practice. As per our implementation, there is an increase of around $25\%$ in training time if $\mathcal{L}_{\text{TR}}{}$ is applied once every 10 $\mathcal{L}_{\text{LM}}{}$ steps, on $25\%$ of attention heads. This excludes the time required for parsing the data used in TreeReg. The parses are either included with the dataset (BLLIP-LG, MultiNLI) or can be computed once and reused for every training run.
@@ 4 Warm-up: Improving Grokking on Sentence Transformation Tasks
We start by training transformer LMs on two diagnostic tasks derived from PCFGs (See Appendix D for examples and data statistics). In Tense Inflection (TI), the model is provided with an input in the past tense, and required to generate the same input in the present tense. In Question Formation (QF), the model is required to transform a declarative sentence into a question. For QF, we report first word accuracy of the decoded question, and for TI, we report the fraction of test inputs for which the target verb is correctly inflected.
@@ Results.
Model Syntactic Generalization ($\uparrow$) PPL ($\downarrow$) Modified Architecture Inference Overhead
BLiMP SG BLLIP PTB
Base LM 72.2 71.9 21.6 49.1 - -
PLM 75.1 80.2 29.8 -
TG - 82.5 30.3 -
Pushdown LM 75.6 82.3 19.9 -
TreeReg LM 74.8 80.0 22.3 44.6
Table 2: Syntactic Generalization on BLiMP and SG test suites, and perplexities (PPL) on BLLIP-LG test set (BLLIP) and PTB test set (PTB) for models trained on BLLIP-LG. Results for PLM, TG and Pushdown LM are taken from Murty et al., 2023b. TreeReg LMs show better syntactic generalization and better generalization to PTB compared to the Base LM.
@@ 5.1 Language Modeling on BLLIP-LG
We train 16-layer LMs on the BLLIP-LG dataset (Hu et al., 2020) (Base LM, detailed hyperparameters in Appendix F). For TreeReg LM, we use $\mathcal{L}_{\text{TR}}{}$ at layer 12 of the LM, on 2 out of 8 attention heads, once every 10 steps of $\mathcal{L}_{\text{LM}}{}$, on the already-parsed BLLIP-LG dataset. A discussion of the layer and attention heads used for $\mathcal{L}_{\text{TR}}{}$ can be found in Appendix G. We benchmark syntactic generalization using the BLiMP and SG test suites (See § 2). To evaluate out-of-distribution generalization, we report perplexity on the Penn TreeBank (PTB; Marcus et al., 1993) test set.
To the best of our knowledge, all prior approaches to induce hierarchical computations in autoregressive causally-trained transformer language models do not leave the underlying transformer architecture untouched, and thus are not directly comparable with TreeReg. Nevertheless, we also report results for Parsing as Language Model (PLMQian et al., 2021), Transformer Grammars (TGSartran et al., 2022), and Pushdown Layers (Pushdown LMMurty et al., 2023b) in Table 2.
@@ Results.
TreeReg achieves a 2.6 point (pt) gain on BLiMP and 8.1 pt overall gain on SG test suites. Figure 2 shows TreeReg improves over the baseline on 4 out of 6 SG test suites, with a notable 17 pt gain on Licensing. TreeReg LM also generalizes better to PTB with 9.2% lower perplexity, with a marginal 0.7 pt perplexity increase on BLLIP-LG.
@@
While PLM, TG and Pushdown LM outperform TreeReg at syntactic generalization on BLiMP and SG test suites, they require additional parameters or incur inference overheads. Except for Pushdown LM, these models also have substantially higher perplexities on the BLLIP-LG test set compared to TreeReg, showing the benefits of the unchanged transformer architecture.
@@
While syntactic generalization improves with more training data for both the Base and TreeReg LMs, TreeReg LMs outperform the baseline across all settings. Moreover, the performance gap widens as the amount of training data increases. Notably, TreeReg surpasses the baseline’s maximum syntactic generalization performance on SG by 4.4 points when trained on only 50% of the data, demonstrating substantially more sample-efficient syntactic generalization. Therefore, even though TreeReg increases training time by 25% compared to a baseline trained for the same number of iterations (see § 3.3), it reaches better syntactic generalization in only 62.5% of the baseline’s training time.
@@ 6.2 Fine-tuning on MultiNLI
We finetune Sheared Llama-1.3B (Base LM) on MultiNLI Williams et al. (2018) (training details and evaluation methodology in Appendix E) to get FT LM. For TreeReg LM, we binarize the parses of the premise and hypothesis sentences provided in the dataset and apply $\mathcal{L}_{\text{TR}}{}$ at layer 16 on 6 out of 24 attention heads, every 10 steps of $\mathcal{L}_{\text{LM}}{}$. We report accuracy on the MultiNLI test set, which includes a matched split from the same sources as the training data, and a mismatched split from different sources. We also present results (Table 5) on two diagnostic datasets: MoNLI Geiger et al. (2020) and MED Yanaka et al. (2019).
@@ 7 Analysis
Although TreeReg only implicitly encodes constituency parses in transformer hidden states, we observe that it demonstrates remarkable parsing capabilities. Parses induced by TreeReg align closely with given parses on all datasets, with F1 scores near or above 90. Parsing performance on PTB and 4000 Questions is slightly better with TreeReg CPT LM, likely due to improved representations from Sheared Llama-1.3B (which benefits from large-scale pre-training) on these datasets, that are relatively out-of-distribution from the BLLIP-LG training data. Notably, induced parses from TreeReg align better with silver parses than Benepar (Kitaev et al., 2022) on the 4000 Questions datasets, by 6.8 F1 pts. We provide some example parses on each dataset in Appendix H.
@@
The F1 scores reveal that at layer 1, the parses are nearly random. As we progress to layer 12, they increasingly align with given parses. Beyond layer 12, there is a sharp drop in F1 scores back to near-random by layer 16. This suggests that the circuit involved in $\mathcal{L}_{\text{TR}}{}$ becomes increasingly aligned with given parses up to layer 12, and then rapidly loses this structure. We conclude that to perform next-word prediction, TreeReg LMs implement a non-syntactic function over the syntactic representations built at layer 12. These syntactic representations are, in turn, built gradually over the preceding layers of the transformer, even in the absence of explicit supervision from TreeReg at these layers.
Training on Randomized Parses.
To assess if the gains from TreeReg are specifically due to hierarchical computation that align closely with constituency parses, we conduct a controlled experiment. We train a 16-layer LM using the exact setup described in § 5.1, but instead of using gold parses, we provide TreeReg with random parses of BLLIP-LG sentences (Randomized TreeReg LM). These randomized parses are generated top-down recursively, by choosing random split points for each sentence span. We then compare performance with the TreeReg LM and Base LM on the BLiMP and SG test suites (Table 7).
Model BLiMP ($\uparrow$) SG ($\uparrow$)
Base LM 72.2 71.9
TreeReg LM 74.8 80
Randomized TreeReg LM 73.4 71.8
Table 7: Results on BLiMP and SG test suites for LMs trained from scratch on BLLIP-LG, with silver and randomized parses (Randomized TreeReg LM) passed to TreeReg. Performance decreases with randomized parses compared to silver parses.
As expected, overall performance on the SG test suites drops marginally compared to Base LM when randomized parses are fed to TreeReg. On the other hand, BLiMP accuracies increase by 1.2 pts even in this setting. We hypothesize that this discrepancy arises from BLiMP’s less granular sentence-level perplexity-based evaluation, in contrast to the surprisal-based approach used in SG.
@@ 8 Other Related Work
A substantial body of research has pointed out the limitations of transformer-based LMs in achieving robust compositional generalization. They have been shown to struggle with simple tasks, such as determining the parity of binary strings (Bhattamishra et al., 2020; Chiang and Cholak, 2022), balancing bracket sequences (Hahn, 2020), as well as more complex diagnostic natural language tasks (McCoy et al., 2020; Geiger et al., 2020; Dziri et al., 2023). Large-scale pre-training does not fully resolve these issues. For instance, Berglund et al. (2024) shows that GPT-4 (Achiam et al., 2023) fails to generalize on bijective relationships (for example, answering “Who is Tom Cruise’s mother?” correctly but not “Who is Mary Lee South’s son?”). Hale and Stanojević, 2024 also find that Gemini Pro (Team et al., 2023) does not learn syntactic universals such as the Final-over-Final condition (Sheehan et al., 2017) for low-resource languages such as Basque.
@@
Building on theories of tree-structured human language processing, prior work explores various tree-structured model architectures (Socher et al., 2013; Tai et al., 2015; Le and Zuidema, 2015; Dyer et al., 2016; Shen et al., 2019; Hu et al., 2021). These models often excel at data-efficient compositional generalization due to their inherently hierarchical computation. However, they have been largely overshadowed by non-hierarchical pre-trained transformer-based LLMs (Brown et al., 2020; Achiam et al., 2023).